# Vec-LUT: Vector Table Lookup for Parallel Ultra-Low-Bit LLM Inference on Edge Devices

### v1

@@ Abstract.
To solve the issue, we propose vector LUT, a new lookup paradigm that constructs a unified LUT across parallel tokens, and performs a single $1\rightarrow N$ lookup per index. To realize it efficiently, we further introduce (1) Vector LUT-Centric Tensor Layout, and (2) Cache-Aware Streamed Lookup techniques. Evaluations on 5 edge devices across 3 LLMs show that Vec-LUT outperforms state-of-the-art baselines by up to $4.2\times$. Our implementation is integrated into llama.cpp. The code is available at https://github.com/Cipherxzc/vlut.cpp.
copyright: nonesubmissionid: 116
@@ 1. Introduction
Large language models (LLMs) are increasingly deployed on edge devices, such as Apple Intelligence on iPhone and Mac (apple2025intelligence), Google Gemma 3 on Android (ikonomidis2025gemma3), and Microsoft Phi on Copilot PC (Mehdi2024; kinfey2024phi3intel). To fit the limited memory budgets of edge devices, quantization has become a de facto technique for on-device LLMs. Recent advances continuously push bit-width from 8-bit (dettmers2022gpt3; xiao2023smoothquant) to 4-bit (frantar2022gptq; lin2023awq), 2-bit (du2024bitdistiller; chen2025efficientqat) and now 1.58-bit, represented by Microsoft’s ternary model and its successors (ma2024158bitllm; ma2025bitnet; 1bitllm2024bitnet; hf1bitllm2024llama3; Falcon3; kaushal2024spectra). These low-bit models not only demonstrate superior bit-level accuracy scaling (huang2025tenet; liu2025paretoq), but also achieve linear speedup via lookup table (LUT)-based inference kernels, enabling CPUs (including Apple M2 and RaspBerry Pi) to outperform NPUs (wei2024tmac; wang2025bitnet). This overturns the need for costly AI accelerators and opens new opportunities for ubiquitous LLM deployment.
@@
LUT-based inference (park2022lut; wei2024tmac) was introduced to bridge the gap between low-bit models and hardware without native low-bit support (ARM2025NEON; Intel2022AVX2; NVIDIA2020AMPERE; Qualcomm2024NPU) (see Fig. 1(b) and Fig. 2). The core idea is to partition the weight matrix into small groups, for which all possible bit patterns of a group can be enumerated. For each bit pattern, its dot product with the activation is precomputed and stored in a lookup table. This turns matrix multiplication into table lookup indexed by the corresponding bit patterns, followed by accumulation across groups. For example, a group of four elements in a ternary model (each weight element ranges {-1, 0, 1}) has 81 $(3^{4})$ possible bit patterns (e.g., (1, -1, 0, 1) and (-1, 0, 1, 0)), resulting in 81 table entries.
@@
Although LUT-based LLM inference achieves superior performance for single-token generation, we find that it behaves poorly under parallel inference (i.e., generating multiple tokens simultaneously), required in scenarios such as prefilling, serving, test-time scaling for decoding (chen2025parallel; ehrlich2025codemonkeys), and speculative decoding (leviathan2023fast; chen2023accelerating). In these settings, the memory bandwidth remains highly underutilized, reaching only $\leq$ 40% of its capacity. This limitation directly harms edge applications that rely on efficient parallel inference, such as multi-app services (shen2025edgelora; li2025mobilora; yuan2024mobile), long-context processing agents (wen2023empowering; shen2025autoiot), and streaming video understanding (ding2025streammind).
@@
Since vector LUT constructs a unified table across all tokens rather than separated ones for each token, all tensors involved in the computation, including weight, activation, and LUT, should be organized to match this paradigm for continuous memory access. In scalar LUT, each token maintains its own table, leading to an activation and weight layout that is feature-first. The unified table used by vector LUT requires a token-first layout so that all tokens’ values for the same feature are stored contiguously. If tensors are misaligned with this requirement, discountinuous memory accesses can degrade performance as high as $12\times$ (see §5.5).
@@
As shown in Fig. 1, vector LUT stores the results for $N$ tokens in each row, which expands the LUT size by a factor of $N$, making it infeasible to cache the entire LUT for fast lookup. For example, while typical L1 and L2/L3 caches for edge processors are only 10s of KiBs and 10s of MiBs, respectively (arm_neoverse_v1; intel_core_i7_13700k), a 2-bit vector LUT in INT16 can exceed $200$ MiB (e.g., $285.5$ MiB for a $4096\times 14436\times 512$ GeMM from Llama3 8B). During lookup, the large LUT combined with the inherently random access pattern can lead to a poor cache hit rate and low inference throughput.
@@
We implement and integrate Vec-LUT into llama.cpp (llamacpp), with particular support for SOTA accuracy-bit-efficient ternary models. Our lossless weight packing can achieve b1.60111The number indicates the BPW (bits per weight) precision (I1) for ternary models, beyond b2.00 (I2), providing the most compact ternary packing to date. We conduct a comprehensive evaluation on 5 edge devices and 3 LLMs (BitNet (1bitllm2024bitnet), Llama 3 (hf1bitllm2024llama3), and Falcon (Falcon3)). Compared to SOTA frameworks (T-MAC (wei2024tmac), bitnet.cpp (ma2025bitnet), and llama.cpp (llamacpp)) requiring only CPUs, Vec-LUT achieves up to $4.2\times$ (I1) and $2.6\times$ (I2) speedup. Besides, Vec-LUT achieves a 273.5 tokens/s throughput on an affordable CPU server (\$0.50/h on AWS) in an additional serving test.
@@
We present Vec-LUT, an efficient mpGeMM kernel built on vector LUT paradigm, optimized with Vector LUT-Centric Tensor Layout and Cache-Aware Streamed Lookup.
@@ 2.1.1. Edge applications of parallel LLM inference
Parallel inference, which processes multiple tokens in parallel, is one of the most critical workloads in today’s LLM-based applications. It is predominantly exercised in two scenarios: prefilling and parallel decoding. (i) Prefilling of LLMs ingests multiple tokens (e.g., in a user prompt) simultaneously as input. Its latency directly affects the responsiveness of edge applications with long context inputs, such as GUI agents (wen2023empowering) and AIoT agents (shen2025autoiot). (ii) Parallel decoding of LLMs generates multiple tokens simultaneously. Its performance not only determines the overall QoS in multi-app services (shen2025edgelora; li2025mobilora), but also influences the effective decoding latency in parallel test-time scaling (ehrlich2025codemonkeys; chen2025parallel) and speculative decoding (leviathan2023fast; chen2023accelerating).
@@ 2.1.2. Superiority of ternary LLMs in edge scenarios
As the bit-width of LLMs continues to shrink, ternary-valued models, where each weight takes value from {1,-1, 0}, emerge as a Pareto-optimal point in the accuracy-cost tradeoff. Their advantages manifest in two dimensions.
(i) Theoretical superiority. Recent research demonstrates the superior accuracy per joule (huang2025tenet) and accuracy per bit trade-offs (liu2025paretoq) of ternary quantization. For example, Microsoft’s BitNet 2B costs only 0.4 GB memory (ma2025bitnet), but 4-bit models of similar accuracy require $\geq$ 1 GB. Also, TENET (huang2025tenet) and LUT Tensor Core (mo2024lut) demonstrate the superior latency/energy efficiency of ternary LLMs with co-designed hardware.
(ii) Practical benefits. Past deployment from Microsoft (e.g., BitNet.cpp and T-MAC) have demonstrated ternary models can run smoothly on diverse x86/ARM/iOS edge devices using CPUs alone, including mobile phones, desktops (Microsoft Surface Book, Apple M2) and even Raspberry Pi. The running speed can even surpass NPUs. For example, T-MAC reports that a 4-bit 7B model reaches 18.7 tokens/s on Snapdragon X Elite CPUs, whereas the NPU attains only 10.4 tokens/s (wei2024tmac). This effectively removes the dependence on GPUs/NPUs for LLM inference and opens significant potential for deployment across edge and IoT devices.
@@ 2.1.3. LUT-based Ultra-Low-Bit LLM inference
LUT-based inference replaces runtime dequantization and multiplication with table lookup. As shown in Fig. 2, take a ternary model with weight partitioned in groups of 4 as an example, it enumerates all possible bit patterns of weight groups (i.e., $\mathbf{w}\in\{-1,0,1\}^{4}$), and then precompute their dot products with the activation group $\mathbf{v}$. The precomputed results are stored in the table $\mathbf{T}=\{\mathbf{w}\times\mathbf{v}\mid\mathbf{w}\in\{-1,0,1\}^{4}\}$, obtaining $3^{4}=81$ table entries in total.
@@ 2.2.1. Costly Dequantization in MAD-based mpGeMM
Conventional MAD-based mpGeMM must dequantize low-bit weights to hardware-supported precisions and then multiply them with activations in aligned precisions (e.g., INT8/FP16). The dequantization and multiplication overheads counteract the benefit of low-bit quantization, and even increase the latency, as demonstrated in Fig. 3222Q# in the figure denotes the quantization formats used in llama.cpp, where # indicates the bits per weight (BPW). TQ# refers to the corresponding ternary quantization formats in llama.cpp.. Among MAD-based mpGeMM kernels in llama.cpp, the higher-BPW 8-bit kernel (Q8_0) delivers lower latency than the 4-bit kernel (Q4_0), and the 2-bit (TQ2_0) ternary packing outperforms the sub-2-bit (TQ1_0) one significantly. In contrast, our LUT-based approach directly leverages compactly packed weights for table lookup without dequantization or unpacking (more in §3.3), yielding remarkable acceleration at low BPWs.
Table 1. Time breakdown of T-MAC’s mpGeMM kernel. Measured with HF BitNet 3B on an Orange Pi 5 Plus.
@@ 2.2.2. Repetitive Lookup in Scalar LUT
Existing LUT-based kernels (e.g., T-MAC) can fully leverage the memory bandwidth for single-token generation. However, we observe a sever bandwidth underutilization (¡40%) for parallel inference. We identify the reason is that existing LUT-based kernels adopt the scalar LUT paradigm ($1\rightarrow 1$ table lookup), which constructs an independent table for each token. Hardware lookup instructions, like SIMD TBL instruction on ARM CPU and PSHUF on x86 CPU is then called to look up the table.
For a parallel inference with $N$ tokens, the kernel requires loading $N$ tables and executing lookup for $N$ times. Since memory access of table lookup is inherently random, and the table size (e.g., hundreds of MiBs) far exceeds cache capacity, this repetitive non-contiguous memory access leads to severe memory-bandwidth underutilization and substantially increased latency. In some cases, scalar-LUT kernels even underperform MAD-based kernels (more in §5.2).
As shown in Table 1, even optimized with hardware instructions, “Lookup” (including weight loading and table lookup) still takes up nearly half of T-MAC’s mpGeMM latency. In contrast, Vec-LUT will reduce the lookup cost to below 1%, as evaluated in §5.4.
@@ 2.3.1. Unified table shared by all tokens
Based on the analysis in §2.2.2, we propose the vector LUT paradigm. Our key insight is that instead of constructing a separate table for each token, we build a unified table shared across all $N$ parallel tokens, replacing $N$ independent $1\rightarrow 1$ lookups with a single $1\rightarrow N$ lookup. This transformation eliminates repeated discontinuous memory accesses, converts them into contiguous accesses for each weight index, and removes the dependency on hardware LUT instructions. Given the high proportion of “Lookup” cost in T-MAC (Table 1), vector LUT is expected to deliver up to $2\times$ kernel-level speedup by effectively eliminating lookup overhead. Vec-LUT reduces lookup cost to below 1% in §5.4.
@@ 2.3.2. Other opportunities by not relying on hardware lookup instructions
Besides the key opportunity in §2.3.1, vector LUT brings two more potential gains by not relying on hardware lookup instructions. (i) Accumulation reduction with larger indexing range. To satisfy the bit-width requirements of SIMD hardware lookup instructions, existing kernels only allow a $2^{4}$ indexing range with a 4-bit index. In contrast, vector LUT allows up to a $3^{5}$ indexing range with an 8-bit index (details in Fig. 5). It reduces the required lookup and accumulation count by $\sim 2\times$ (twice lookup and accumulation for 4-bit index compared to 8-bit index), providing up to $1.25\times$ additional kernel speedup (see Table 1 accumulate cost). (ii) Enhanced versatility and efficiency with flexible sub-2-bit packing. To utilize SIMD acceleration, existing MAD-based kernels impose strict weight shape requirements (e.g., by multiples of 256 in llama.cpp’s TQ1_0 and TQ2_0 (compilade2024tq)), which significantly limits their generality. In contrast, vector LUT can pack weights of almost any shapes more compactly for memory saving, without compromising performance or accuracy.

### v2

@@ Abstract.
To solve the issue, we propose vector LUT, a new lookup paradigm that constructs a unified LUT across parallel tokens, and performs a single $1\rightarrow N$ lookup per index. To realize it efficiently, we further introduce (1) Vector LUT-Centric Tensor Layout, and (2) Cache-Aware Streamed Lookup techniques. Evaluations on 5 edge devices across 3 LLMs show that Vec-LUT outperforms state-of-the-art baselines by up to $4.2\times$. Our implementation is integrated into llama.cpp. The code is available at https://github.com/OpenBitSys/vlut.cpp.
Large Language Models, Low-Bit Quantization, General Matrix Multiplication
journalyear: 2026copyright: ccconference: The 24th Annual International Conference on Mobile Systems, Applications and Services; June 21–25, 2026; Cambridge, United Kingdombooktitle: The 24th Annual International Conference on Mobile Systems, Applications and Services (MobiSys ’26), June 21–25, 2026, Cambridge, United Kingdomdoi: 10.1145/3745756.3809200isbn: 979-8-4007-2027-7/2026/06ccs: Computing methodologies Parallel algorithmsccs: Human-centered computing Ubiquitous and mobile computing
@@ 1. Introduction
Figure 2. How mpGeMM kernel optimization integrates into the high-level Transformer inference flow. In a quantized LLM, the linear operators, which dominate inference latency in many scenarios, perform mixed-precision multiplication between weights (e.g., ternary) and activations (e.g., INT8/FP16). Our method optimizes memory access within the mpGeMM kernel, thereby improving overall inference performance and efficiency.
@@
Large language models (LLMs) are increasingly deployed on edge devices, such as Apple Intelligence on iPhone and Mac (Apple Inc., 2025), Google Gemma 3 on Android (Ikonomidis et al., 2025), and Microsoft Phi on Copilot PC (Mehdi, 2024; kinfey, 2024). To fit the limited memory budgets of edge devices, quantization has become a de facto technique for on-device LLMs. Recent advances continuously push bit-width from 8-bit (Dettmers et al., 2022; Xiao et al., 2023) to 4-bit (Frantar et al., 2023; Lin et al., 2024), 2-bit (Du et al., 2024; Chen et al., 2025a) and now 1.58-bit, represented by Microsoft’s ternary model and its successors (Ma et al., 2024, 2025; 1bitLLM, 2024; Mekkouri et al., 2024; Falcon-LLM Team, 2024; Kaushal et al., 2025). These low-bit models not only demonstrate superior bit-level accuracy scaling (Huang et al., 2025; Liu et al., 2025), but also achieve linear speedup via lookup table (LUT)-based inference kernels, enabling CPUs (including Apple M2 and Raspberry Pi) to outperform NPUs (Wei et al., 2025; Wang et al., 2025). This overturns the need for costly AI accelerators and opens new opportunities for ubiquitous LLM deployment.
@@
LUT-based inference (Park et al., 2024; Wei et al., 2025) was introduced to bridge the gap between low-bit models and hardware without native low-bit support (Arm Limited, 2025a; Intel Corporation, 2022b; NVIDIA Corporation, 2020; Qualcomm Technologies, Inc., 2024) (see Fig. 1(b), Fig. 2, and Fig. 3). The core idea is to partition the weight matrix into small groups, for which all possible bit patterns of a group can be enumerated. For each bit pattern, its dot product with the activation is precomputed and stored in a lookup table. This turns matrix multiplication into table lookup indexed by the corresponding bit patterns, followed by accumulation across groups. For example, a group of four elements in a ternary model (each weight element ranges {-1, 0, 1}) has 81 $(3^{4})$ possible bit patterns (e.g., (1, -1, 0, 1) and (-1, 0, 1, 0)), resulting in 81 table entries.
@@
Although LUT-based LLM inference achieves superior performance for single-token generation, we find that it behaves poorly under parallel inference (i.e., generating multiple tokens simultaneously), required in scenarios such as prefilling, serving, test-time scaling for decoding (Chen et al., 2025b; Ehrlich et al., 2025), and speculative decoding (Leviathan et al., 2023; Chen et al., 2023). In these settings, the memory bandwidth remains highly underutilized, reaching only $\leq$ 40% of its capacity. This limitation directly harms edge applications that rely on efficient parallel inference, such as multi-app services (Shen et al., 2025b; Li et al., 2025; Yuan et al., 2024), long-context processing agents (Wen et al., 2024; Shen et al., 2025a), and streaming video understanding (Ding et al., 2025).
@@
Since vector LUT constructs a unified table across all tokens rather than separated ones for each token, all tensors involved in the computation, including weight, activation, and LUT, should be organized to match this paradigm for contiguous memory access. In scalar LUT, each token maintains its own table, leading to an activation and weight layout that is feature-first. The unified table used by vector LUT requires a token-first layout so that all tokens’ values for the same feature are stored contiguously. If tensors are misaligned with this requirement, non-contiguous memory accesses can degrade performance by up to $12\times$ (see §5.5).
@@
As shown in Fig. 1, vector LUT stores the results for $N$ tokens in each row, which expands the LUT size by a factor of $N$, making it infeasible to cache the entire LUT for fast lookup. For example, while typical L1 and L2/L3 caches for edge processors are only 10s of KiBs and 10s of MiBs, respectively (Arm Limited, 2025b; Intel Corporation, 2022a), a 2-bit vector LUT in INT16 can exceed $200$ MiB (e.g., $285.5$ MiB for a $4096\times 14436\times 512$ GeMM from Llama3 8B). During lookup, the large LUT combined with the inherently random access pattern can lead to a poor cache hit rate and low inference throughput.
@@
[…]
@@
We present Vec-LUT, an efficient mpGeMM kernel built on the vector LUT paradigm, optimized with Vector LUT-Centric Tensor Layout and Cache-Aware Streamed Lookup.
@@ 2.1.1. Edge applications of parallel LLM inference
Parallel inference, which processes multiple tokens in parallel, is one of the most critical workloads in today’s LLM-based applications. It is predominantly exercised in two scenarios: prefilling and parallel decoding. (i) Prefilling of LLMs ingests multiple tokens (e.g., in a user prompt) simultaneously as input. Its latency directly affects the responsiveness of edge applications with long context inputs, such as GUI agents (Wen et al., 2024) and AIoT agents (Shen et al., 2025a). (ii) Parallel decoding of LLMs generates multiple tokens simultaneously. Its performance not only determines the overall QoS in multi-app services (Shen et al., 2025b; Li et al., 2025), but also influences the effective decoding latency in parallel test-time scaling (Ehrlich et al., 2025; Chen et al., 2025b) and speculative decoding (Leviathan et al., 2023; Chen et al., 2023). Accelerating parallel inference on edge devices is therefore essential to meet the latency and throughput requirements of these applications, enabling the deployment of more capable models under resource constraints.
@@ 2.1.2. Superiority of ternary LLMs in edge scenarios
As the bit-width of LLMs continues to shrink, ternary-valued models, where each weight takes value from {1,-1, 0}, emerge as a Pareto-optimal point in the accuracy-cost tradeoff. Their advantages manifest in two dimensions. (i) Theoretical superiority. Recent research demonstrates the superior accuracy per joule (Huang et al., 2025) and accuracy per bit trade-offs (Liu et al., 2025) of ternary quantization. For example, Microsoft’s BitNet 2B costs only 0.4 GB memory (Ma et al., 2025), but 4-bit models of similar accuracy require $\geq$ 1 GB. Also, TENET (Huang et al., 2025) and LUT Tensor Core (Mo et al., 2025) demonstrate the superior latency/energy efficiency of ternary LLMs with co-designed hardware. (ii) Practical benefits. Past deployments from Microsoft (e.g., BitNet.cpp and T-MAC) have demonstrated ternary models can run smoothly on diverse x86/ARM/iOS edge devices using CPUs alone, including mobile phones, desktops (Microsoft Surface Book, Apple M2) and even Raspberry Pi. The running speed can even surpass NPUs. For example, T-MAC reports that a 4-bit 7B model reaches 18.7 tokens/s on Snapdragon X Elite CPUs, whereas the NPU attains only 10.4 tokens/s (Wei et al., 2025). This effectively removes the dependence on GPUs/NPUs for LLM inference and opens significant potential for deployment across edge and IoT devices.
@@ 2.1.3. LUT-based Ultra-Low-Bit LLM inference
LUT-based inference replaces runtime dequantization and multiplication with table lookup. As shown in Fig. 3, take a ternary model with weights partitioned in groups of 4 as an example, it enumerates all possible bit patterns of weight groups (i.e., $\mathbf{w}\in\{-1,0,1\}^{4}$), and then precomputes their dot products with the activation group $\mathbf{v}$. The precomputed results are stored in the table $\mathbf{T}=\{\mathbf{w}\times\mathbf{v}\mid\mathbf{w}\in\{-1,0,1\}^{4}\}$, obtaining $3^{4}=81$ table entries in total.
@@ 2.2.1. Costly Dequantization in MAD-based mpGeMM
Conventional MAD-based mpGeMM must dequantize low-bit weights to hardware-supported precisions and then multiply them with activations in aligned precisions (e.g., INT8/FP16). The dequantization and multiplication overheads counteract the benefit of low-bit quantization, and even increase the latency, as demonstrated in Fig. 4222Q# in the figure denotes the quantization formats used in llama.cpp, where # indicates the bits per weight (BPW). TQ# refers to the corresponding ternary quantization formats in llama.cpp.. Among MAD-based mpGeMM kernels in llama.cpp, the higher-BPW 8-bit kernel (Q8_0, dequantization-free) delivers lower latency than the 4-bit kernel (Q4_0), and the 2-bit (TQ2_0) ternary packing outperforms the sub-2-bit (TQ1_0) one significantly. In contrast, our LUT-based approach directly leverages compactly packed weights for table lookup without dequantization or unpacking (more in §3.3), yielding remarkable acceleration at low BPWs.
Table 1. Time breakdown of T-MAC’s mpGeMM kernel. Measured on an Orange Pi 5 Plus with a single thread.
@@ 2.2.2. Repetitive Lookup in Scalar LUT
Existing LUT-based kernels (e.g., T-MAC) can fully leverage the memory bandwidth for single-token generation. However, we observe a severe bandwidth underutilization (¡40%) for parallel inference. We identify the reason is that existing LUT-based kernels adopt the scalar LUT paradigm ($1\rightarrow 1$ table lookup), which constructs an independent table for each token, and then calls hardware lookup instructions (e.g., SIMD TBL instruction on ARM CPU and PSHUF on x86 CPU) to look up the table.We find that this paradigm fundamentally cannot achieve efficient reuse of both tables and indices simultaneously.
During parallel inference with $N$ tokens, the kernel must construct $N$ tables, each randomly accessed $\mathcal{O}(MK)$ times. T-MAC reuses each table across all indices before loading the next, but incurs $N\times$ repetitive table loading and lookup with non-contiguous memory access within tables. The alternative—reusing each index across tables—would cause non-contiguous memory access across tables or even frequent LUT swapping, with even higher overhead. Both reuse strategies in scalar LUT lead to severe memory-bandwidth underutilization and substantially increased latency. In some cases, scalar-LUT kernels even underperform MAD-based kernels (more in §5.2).
As shown in Table 1, even optimized with hardware instructions, “Lookup” (including weight loading and table lookup) still takes up nearly half of T-MAC’s single-thread mpGeMM latency. In contrast, Vec-LUT will reduce the lookup cost to below 1% (details in §5.4).
@@ 2.3.1. Unified table shared by all tokens
Based on the analysis in §2.2.2, we propose the vector LUT paradigm. Our key insight is that instead of constructing a separate table for each token, we build a unified table shared across all $N$ parallel tokens, replacing $N$ independent $1\rightarrow 1$ lookups with a single $1\rightarrow N$ lookup. This transformation eliminates repeated non-contiguous memory accesses, converts them into contiguous accesses for each weight index, and removes the dependency on hardware LUT instructions. Given the high proportion of “Lookup” cost in T-MAC (Table 1), vector LUT is expected to deliver up to $2\times$ kernel-level speedup by effectively eliminating lookup overhead. Vec-LUT reduces lookup cost to below 1% in §5.4.
@@ 2.3.2. Other opportunities by not relying on hardware lookup instructions
Besides the key opportunity in §2.3.1, vector LUT brings two more potential gains by not relying on hardware lookup instructions. (i) Accumulation reduction with larger indexing range. To satisfy the bit-width requirements of SIMD hardware lookup instructions, existing kernels only allow a $2^{4}$ indexing range with a 4-bit index. In contrast, vector LUT allows up to a $3^{5}$ indexing range with an 8-bit index (details in Fig. 6). It reduces the required lookup and accumulation count by $\sim 2\times$ (twice lookup and accumulation for 4-bit index compared to 8-bit index), providing up to $1.25\times$ additional kernel speedup (see Table 1 accumulate cost). (ii) Enhanced versatility and efficiency with flexible sub-2-bit packing. To utilize SIMD acceleration, existing MAD-based kernels impose strict weight shape requirements (e.g., by multiples of 256 in llama.cpp’s TQ1_0 and TQ2_0 (compilade, 2024)), which significantly limits their generality. In contrast, vector LUT can pack weights of almost any shapes more compactly for memory saving, without compromising performance or accuracy.